# Some classes satisfying the 2-dimensional Jacobian conjecture

### v3

@@ 3. Two-dimensional complex Jacobian conjecture of degree until 104
In this case ${\rm deg}(u_{0})=1$, ${\rm deg}(u_{1})=3$ and ${\rm deg}(u_{3})=6$. Then if ${\rm deg}(F)=105$, it is possible to reduce ${\rm deg}(F)$ to only 13, choosing for example $u_{2}u_{3}^{12}$ as a term of the non-proper map $F$. More precisely, each polynomial $F_{1}$ and $F_{2}$ is a finite sum of monomials $u_{0}^{\alpha_{0}}u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}$ such that $\alpha_{0}+3\alpha_{1}+6\alpha_{2}\leq 105$. Besides, with a fixed number $n$ of variables, then by increasing the natural numbers $r$ and $s$, we can reduce quickly the degree of the map.
Statements and Declarations
The author declares that she has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper. The author was supported by the Research Support Foundation of Sao Paulo State (FAPESP), grants 2023/07802-7 and 2019/21181-0.

### v5

@@ 3. Two-dimensional complex Jacobian conjecture of degree until 104
In this case ${\rm deg}(u_{0})=1$, ${\rm deg}(u_{1})=3$ and ${\rm deg}(u_{2})=6$. Then if ${\rm deg}(F)=105$, each polynomial $F_{1}$ and $F_{2}$ is a finite sum of monomials $u_{0}^{\alpha_{0}}u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}$ such that $\alpha_{0}+3\alpha_{1}+6\alpha_{2}\leq 105$.
Moreover, with a fixed number $n$ of variables, then by increasing the natural numbers $r$ and $s$, we can reduce quickly the degree of the map.
Remark 3.9.
One can completely generalize the definition of “non-proper variables” (Definition 2.2) for the general case $F:\mathbb{K}^{n}\to\mathbb{K}^{n}$, $n\geq 3$, for both real and complex cases $\mathbb{K}=\mathbb{R}$ and $\mathbb{K}=\mathbb{C}$.
Acknowledgement.
The author would like to thank Prof. Luis Renato Gonçalves Dias for many useful discussions during her visit to the Universidade Federal de Uberlândia (Brazil) in December of 2019.
This paper was financially supported by the Brazilian Research Support Foundation of Sao Paulo State (FAPESP) process 2023/07802-7 and by the Brazilian National Council for Scientific and Technological Development project CNPq ${\rm n}^{\rm o}$:407454/2023-3.